# LabObf: A Label Protection Scheme for Vertical Federated Learning Through Label Obfuscation

### v1

@@ Abstract
Split learning, as one of the most common architectures in vertical federated learning, has gained widespread use in industry due to its privacy-preserving characteristics. In this architecture, the party holding the labels seeks cooperation from other parties to improve model performance due to insufficient feature data. Each of these participants has a self-defined bottom model to learn hidden representations from its own feature data and uploads the embedding vectors to the top model held by the label holder for final predictions. This design allows participants to conduct joint training without directly exchanging data. However, existing research points out that malicious participants may still infer label information from the uploaded embeddings, leading to privacy leakage. In this paper, we first propose an embedding extension attack that manually modifies embeddings to undermine existing defense strategies, which rely on constraining the correlation between the embeddings uploaded by participants and the labels. Subsequently, we propose a new label obfuscation defense strategy, called ‘LabObf’, which randomly maps each original one-hot vector label to multiple numerical soft labels with values intertwined, significantly increasing the difficulty for attackers to infer the labels. We conduct experiments on four different types of datasets, and the results show that LabObf can reduce the attacker’s success rate to near random guessing while maintaining an acceptable model accuracy.
@@ 1 Introduction
In the era of data deluge, the resources held by individuals or individual entities, whether in terms of hardware or software, are gradually proving inadequate to handle the exponentially growing volume of data and the extensive task of mining potential user demands. Vertical federated learning (VFL) has emerged as a privacy-preserving collaborative framework in response to this challenge. Compared to traditional horizontal federated learning, VFL is more suitable for scenarios where participants hold data with different attribute features and cannot directly share them. It finds extensive applications in fields such as healthcare[1], finance[2], and the internet of things (IoT)[3].
@@
Split learning[4] is one of the most widely used organizational structures in real-world VFL scenarios. In this setup, the dataset is vertically divided into several parts, each held by different participants (usually two parties), with label information held exclusively by one participant. The participant holding the labels is typically called the label party or the host, while the other participants are referred to as clients. The host is typically the initiator of the federated task, seeking to improve the primary task’s performance by aggregating data from other entities (i.e., clients). Generally, the host has both a bottom model and a top model, while each client holds its own bottom model. The participants use their bottom models to extract embeddings from their respective feature dimensions and then upload these embeddings to the host. The host concatenates the embeddings into a single input for the top model, which then outputs the final prediction. The layer used for interaction between the bottom models and the top model is called the cut layer. This structure is adopted by several renowned companies, such as ByteDance[5] and Tencent[6], for two-party VFL.
@@
Although such architecture has been proven to possess strong privacy protection capabilities, existing research indicates that they still suffer from security vulnerabilities[7, 8, 9, 10]. Data owned by participants, especially label data crucial for prediction tasks, is at risk of being stolen. These risks often occur during the information exchange process at the cut layer. For instance, malicious clients may steal the labels by exploiting the back-propagation of gradients from the top model[7, 8] or the distribution of embeddings in the cut layer[9]. The former methods, such as the shortest path attack[7], entails inferring label information by assessing the distance between back-propagated gradients of target samples and those from the 00-class and $1$-class samples. However, these attacks are primarily utilized in binary classification tasks and there are relatively mature solutions available to defend against them[7, 8]. The latter methods, represented by the label inference attack[9], capitalizes on the strong correlation between the embedding vectors from the local bottom model and labels. Attackers try to gather a small set of labeled samples to mimic training the top model through semi-supervised learning. Using the local ‘top model’, they can then deduce label information from the intermediate data transmitted through the cut layer with high confidence. This method is highly adaptable, poses a significant threat to VFL models, and is currently a prevalent attack strategy. Moreover, in real-world scenarios, clients often possess the majority, or even all, of the attributes. In such cases, this attack would be more effective and exhibit greater aggressiveness.
@@
In response to the second threat of label leakage, some works attempt to protect labels through perturbation [11, 12, 13]. These methods disrupt attacks that aim to extract label information from embeddings by carefully designing fine-grained noise applied to back-propagated gradients or embeddings. However, introducing random noises into these critical intermediate results will inevitably impact the training accuracy of the main task. Therefore, to achieve a good balance between security and model accuracy, the magnitude of the noises must be limited. As a result, when attackers are highly capable, e.g, with a larger auxiliary set of labeled samples, these schemes tend to lack robust defense.
@@
Another representative defense approach is Discorloss[14], which introduces an additional term in the loss function to constrain the correlation between the client’s local embeddings and the labels, thereby increasing the difficulty for malicious clients to infer labels. However, this method primarily suppresses the direct distance correlation between the client’s embeddings and labels, which might be disrupted by the malicious client as the corresponding embeddings are fully under her/his control. This paper introduces such an attack called the embedding extension attack, which actively reduces the overall correlation between the local embeddings and labels by introducing some useless perturbation dimensions into the embedding vectors, thereby disabling the defense mechanism.
@@
To address the security vulnerabilities faced by current vertical federated learning, such as split neural networks[4] (SplitNN), this paper proposes a label obfuscation defense mechanism called LabObf, which is highly effective and less susceptible to attacker interference, with the aim of specifically safeguarding label data from leakage. In this scheme, the host, as the defending party, secretly remaps the original labels to a new distribution space in a one-to-multiple manner. For instance, in the case of a binary classification dataset, the original 00-class and $1$-class labels are each mapped to four real-valued soft labels. These eight soft labels are then scattered and exhibit a cross-distribution state in their values (for example, $0\to[0.1,0.3,0.6,0.9]$, $1\to[0.2,0.4,0.5,0.8]$). It is important to note that this process occurs on the host side, unknown to the client, and for each original label class, the mapping to each soft label is uniform random. This design aims to obscure the distribution of original class features at the cut layer through label obfuscation, making it harder for attackers to infer label information from their local embeddings. However, this approach presents a challenge in that training the main task of the model becomes more difficult, as the original data features lack a direct mapping to the soft labels. To ensure that the model learns this artificial mapping, modifications need to be made to the dataset, such as artificially constructing additional features to establish the mapping from feature data to soft labels. During this process, the host and client each hold a portion of additional features, and both parts jointly determine the mapping relationship to prevent the client from easily modifying or inferring the mapping.
@@
By training with the additional features and soft labels, the approach adds a layer of protection to reduce the direct information about real labels that could be inferred from the model parameters. The mapping between original labels and soft labels is controlled by the label holder, making it challenging for attackers to access or disrupt this process. Furthermore, this scheme partially mitigates the conflict between primary task accuracy and defense strength that perturbation-based defense approaches often face. In this approach, the higher the training accuracy with soft labels, the higher the accuracy of the converted real labels, without compromising defense performance. This is because the defense strength is determined by the degree of obfuscation in the mapping relationship of the soft labels. Finally, such a scheme is practical and straightforward to implement in real-world scenarios.
@@
  • Validate the security of existing defense works: We propose a disruption method that renders current defense solutions ineffective against the label inference attack[9].
  • Propose a robust defense mechanism against interference: We map original labels to a newly confused distribution, reducing the risk of label information leakage at the cut layer.
  • Conduct detailed empirical validation: We conduct experiments on several popular datasets including Epsilon[15], Bank[16], CovType[17], and Fault Type[18]. Among these, Epsilon and Bank are binary classification datasets, while CovType and Fault Type are multi-class classification datasets. The experimental results demonstrate that our proposed approach exhibits strong defense capabilities. The success rate of attacks decreased by up to $16\%$, approaching the level of random guessing. Furthermore, the model accuracy decreased by less than $6\%$, maintaining an acceptable range.
In the following sections, we describe VFL and the existing background knowledge of attacks and defenses in Sec.2. The threat model is presented in Sec.3. We first present an attack to demonstrate the flaw in the existing defense scheme in Sec.4, and then propose a new, more robust protection scheme in Sec.5. Sec.6 presents experimental data and analysis, while Sec.8 analyzes related works. Finally, we conclude with a brief summary in Sec.9.
@@ 2.1 SplitNN-Based VFL
Although SplitNN can be applied to settings with multiple parties, in practice, VFL is often limited to two parties[7, 9, 20, 11, 14, 21] due to issues of trust and efficiency. Therefore, this paper focuses exclusively on scenarios with two-party participation.
@@ 3.1 Network Architecture
Due to the widespread application of VFL tasks in practical environments involving two-party participation[7, 9, 20, 11, 14, 21], this paper primarily focuses on the SplitNN architecture involving two participating parties, namely the host and the client, like Fig.1. In this architecture, the client, potentially acting as an attacker, holds partial dimensional features and maintains the corresponding bottom model locally. During training and prediction, it is responsible for uploading locally computed embeddings to the host. The host, acting as the defending party, holds partial dimensional features and all labels, and maintains both the corresponding bottom and top models locally. During training and prediction, it calculates the corresponding embeddings locally and combines them with the embeddings uploaded by the client, and then inputs them into the top model. The output of the top model represents the prediction result.
@@ 4 Embedding Extension Attack against Discorloss
Figure 3: The Pearson correlation between the embedding dimensions and the real labels.
@@ 5 Label Obfuscation Protection Scheme
(a) Raw model.
(b) The client has no additional attributes.
(c) The client has an additional attribute.
Figure 7: An example of how to build the mapping relationship between real classes and fake classes using the additional attributes.
@@ 6.3 Evaluation of LabObf
We validate the performance of LabObf on four datasets. In the experiments, we add one-dimensional special features, $x_{1}$ and $x_{2}$, to the attacker and the defender respectively. These features have values randomly selected from the interval $[0,200]$. Utilizing these two special features, we map each original label $l$ to a pair of soft labels $[l_{1},l_{2}]$. The specific mapping rule is as follows: when $x_{1}+x_{2}\leq 200$, $l$ is mapped to $l_{1}$; otherwise, $l$ is mapped to $l_{2}$. The mapping relationship is designed in a way that the soft label values from the same original label are as far apart as possible, while the soft label values from different original labels are as close as possible. The specific mapping relationships are as follows: For Epsilon, we map label 00 to soft labels $[0,0.8]$, and label $1$ to soft labels $[0.2,1]$. For Bank, we map label 00 to soft labels $[0,0.67]$, and label $1$ to soft labels $[0.33,1]$. For CovType, we map the seven original labels to the following soft labels respectively: $[0.0,3.5]$, $[0.5,4.0]$, $[1.0,4.5]$, $[1.5,5.0]$, $[2.0,5.5]$, $[2.5,6.0]$, $[3.0,6.5]$. For Fault, we map the three original labels to the following soft labels respectively: $[0,1.5]$, $[0.5,2]$, $[1,2.5]$. We train VFL models based on these soft labels and then applied label inference attacks to exploit predictions from the perturbed bottom model outputs.
@@ 7.1 Effect of Auxiliary Set Size on LabObf
Figure 12: Model performance and attack performance on binary classification tasks under different selection of soft labels.
7.2 Effect of Soft Label Selection on LabObf
The selection of soft labels significantly impacts the effectiveness of LabObf. This can be explained in three aspects: the magnitude of soft labels, the degree of cross-over between soft labels, and the number of soft labels. We analyze these three aspects and provide design principles as follows:
  • The magnitude of soft labels is significant. When decomposing the original label into soft labels, larger soft label values lead to larger gradients within the loss function, consequently slowing down the model convergence. Additionally, larger label values may introduce numerical instability during training. Conversely, decreasing the label values might cause the model to overlook subtle differences in the data, potentially leading to the model’s inability to accurately capture important patterns and features, thus reducing its predictive capability. Therefore, we choose to normalize the data range of soft labels to match the original label data range, mitigating additional training difficulties caused by label decomposition.
  • The degree of cross-over between soft labels provides a means to adjust the defense strength. If the soft labels are identical to the original label, the degree of cross-over will be minimized, and our defense scheme will degrade to no defense at all. Conversely, if the soft labels are interleaved pairwise, our defense effectiveness will be maximized. Hence, the degree of cross-over reflects the defense strength, allowing us to adjust it for a better privacy-utility trade-off.
  • The number of soft labels into which each original label is split also plays a crucial role in the defense’s effectiveness. Having too many soft labels may increase the learning task’s difficulty, requiring more complex models and additional iterations, while having too few soft labels may fail to achieve deep cross-interactions. Additionally, the number of soft labels affects the complexity and computational overhead of the defense scheme. Generally, splitting each original label into $2$ to $4$ soft labels is more appropriate.
Next, we present the performance of LabObf under different soft label selections that follow these design principles. To simplify the analysis, we assume each original label maps to an equal number of soft labels, considering three settings where each original label maps to $2$, $3$, or $4$ soft labels. Each setting is optimized according to the design principles for soft label magnitude and cross-over degree. Given that the same settings on multiple classification datasets would generate too many soft labels, we experiment on the binary classification datasets Epsilon and Bank.
Table 12 illustrates the model performance and attack performance under different selections of soft labels. We observe that as the number of soft labels increases, the primary task accuracy decreases, and the attack success rate increases. This is reasonable, as an increased number of soft labels makes it harder for the model to learn and fit the soft label distribution, leading to a decrease in soft label classification accuracy, which in turn lowers the original label classification accuracy and defense strength. It is important to note that even so, the model accuracy and attack success rates remain within an acceptable range. This indicates that by following the design principles mentioned above, LabObf can achieve a satisfactory defense performance for label protection.

### v2

@@ Abstract
Split Neural Network, as one of the most common architectures used in vertical federated learning, is popular in industry due to its privacy-preserving characteristics. In this architecture, the party holding the labels seeks cooperation from other parties to improve model performance due to insufficient feature data. Each of these participants has a self-defined bottom model to learn hidden representations from its own feature data and uploads the embedding vectors to the top model held by the label holder for final predictions. This design allows participants to conduct joint training without directly exchanging data. However, existing research points out that malicious participants may still infer label information from the uploaded embeddings, leading to privacy leakage. In this paper, we first propose an embedding extension attack manipulating embeddings to undermine existing defense strategies, which rely on constraining the correlation between the embeddings uploaded by participants and the labels. Subsequently, we propose a new label obfuscation defense strategy, called ‘LabObf’, which randomly maps each original integer-valued label to multiple real-valued soft labels with values intertwined, significantly increasing the difficulty for attackers to infer the labels. We conduct experiments on four different types of datasets, and the results show that LabObf significantly reduces the attacker’s success rate compared to raw models while maintaining desirable model accuracy.
@@ 1 Introduction
In the era of data deluge, the resources held by individuals or agencies, whether in terms of hardware or software, are gradually proved to be inadequate to handle the exponentially growing volume of data and the extensive task of mining potential user demands. Vertical federated learning (VFL) has emerged as a multi-party collaborative framework in response to this challenge. Compared to traditional horizontal federated learning, VFL is more suitable for scenarios where participants hold data with different features and do not want to share them directly. It finds extensive applications in fields such as healthcare[1], finance[2], and Internet of Things (IoT)[3].
@@
Split Neural Network[4] is one of the most widely used organizational structures in real-world VFL scenarios. In this setup, the dataset is vertically divided into several parts, each held by different participants (usually two parties), with label information held exclusively by one participant. The participant holding the labels is typically called the label party or the host, while the other participants are referred to as clients. The host is typically the initiator of the federated task, seeking to improve the primary task’s performance by aggregating data from other entities (i.e., clients). Generally, the host has both a bottom model and a top model, while each client holds its own bottom model. The participants use their bottom models to learn embeddings from their own feature data and then upload these embeddings to the host. The host concatenates the embeddings into a single input for the top model, which then outputs the final prediction. The layer used for interaction between the bottom models and the top model is called the cut layer. This structure is adopted by several renowned companies and communities, such as ByteDance[5], Tencent[6], and OpenMined[7], for two-party VFL.
@@
Although such architecture has been proven to possess strong privacy protection capabilities[8, 9], existing research indicates that they still suffer from security vulnerabilities[10, 11, 12, 13]. Data owned by participants, especially label data crucial for prediction tasks, is at risk of being stolen. These risks exist in the information exchange process at the cut layer. For instance, malicious clients may steal the labels by exploiting the back-propagation of gradients from the top model[10, 11] or the distribution of embeddings in the cut layer[12]. The former methods, such as the shortest path attack[10], entails inferring label information by assessing the distance between back-propagated gradients of target samples and those from the 00-class and $1$-class samples. However, these attacks are primarily utilized in binary classification tasks and there are relatively mature solutions available to defend against them[10, 11]. Therefore, this type of attack will not be discussed further in this paper.
The latter methods, which is also our main consideration, represented by the label inference attack[12], capitalizes on the strong correlation between the embedding vectors from the local bottom model and labels. Attackers try to gather a small set of labeled samples to mimic training the top model through semi-supervised learning. Using the local ‘top model’, they can then deduce label information from the intermediate data transmitted through the cut layer with high confidence. This method is highly adaptable, poses a significant threat to VFL models, and is currently a prevalent attack strategy. Moreover, in real-world scenarios, clients often possess the majority, or even all, of the attributes. In such cases, this attack would be more effective and exhibit greater aggressiveness.
@@
In response to this kind of label leakage, some works attempt to protect labels through perturbation [14, 15, 16]. These methods essentially add fine-grained noise to the embeddings or back-propagated gradients to blur the differences between embeddings of different classes without affecting the main task’s accuracy, thereby misleading or disrupting label inference attacks. This approach is advantageous because it is lightweight and does not require modifications to the main task, yielding good results when the attacker’s capabilities are limited (e.g., with an auxiliary set of fewer than $100$ samples). However, such noise-adding schemes inevitably creates conflicts with the main task training due to its underlying principles. Therefore, to balance data security and model accuracy, the magnitude of the noise must be limited. Consequently, when attackers have higher capabilities, such as a slightly larger auxiliary set of labeled samples (e.g., $500$ samples, which is still within the attacker’s capability but greatly exceeds the scope considered by such schemes), these schemes tend to lack robust defense.
@@
Another representative defense approach is Discorloss[17], which introduces an additional term in the loss function to constrain the correlation between the client’s local embeddings and the labels, thereby increasing the difficulty for malicious clients to infer labels. This approach leverages the inherent redundancy of neural networks to help find a balance between main task accuracy and defense performance. However, this method primarily suppresses the direct distance correlation between the client’s embeddings and labels, which might be disrupted by the malicious client as the corresponding embeddings are fully under her/his control.
To further expose the above security vulnerabilities, we first introduce an attack called the Embedding Extension Attack to validate the flaw of Discorloss, demonstrating that it is risky to deploy defense by leveraging the client. Given that Discorloss relies on embeddings uploaded by the client for its calculations, the malicious client can add several perturbation dimensions to its output embedding. By assigning carefully crafted special values to these perturbation dimensions, the client can maintain a high correlation between the original embedding and the label while reducing the correlation between the embedding uploaded to the host and the label. This undermines Discorloss’s constraint on the original embedding. Thus, as an attacker, it can still use the original embeddings that have evaded the constraints to infer labels.
@@
Next, we propose a label obfuscation defense mechanism called LabObf, which is highly effective and less susceptible to attacker interference, with the aim of specifically safeguarding label data from leakage. In this scheme, the host, as the defending party, secretly remaps the original labels to a new distribution space in a one-to-multiple manner. For instance, in the case of a binary classification dataset, the original 00-class and $1$-class labels are each mapped to four real-valued soft labels. These eight soft labels are then scattered and exhibit a cross-distribution state in their values (for example, $0\to[0.1,0.3,0.6,0.9]$, $1\to[0.2,0.4,0.5,0.8]$). It is important to note that this process occurs on the host side, unknown to the client, and for each original label class, the mapping to each soft label is uniform random. This design aims to obscure the distribution of original class features at the cut layer through label obfuscation, making it harder for attackers to infer label information from their local embeddings. However, this approach presents a challenge in that training the main task of the model becomes more difficult, as the original data features lack a direct mapping to the soft labels. To ensure that the model learns this artificial mapping, modifications need to be made to the dataset, such as artificially constructing additional features to establish the mapping from feature data to soft labels. During this process, the host and client each hold a portion of additional features, and both parts jointly determine the mapping relationship to prevent the client from easily modifying or inferring the mapping.
@@
LabObf addresses some of the shortcomings of previous defense methods. On the one hand, this scheme partially mitigates the conflict between primary task accuracy and defense strength that perturbation-based defense approaches often face. In this approach, the higher the training accuracy with soft labels, the higher the accuracy of the converted real labels, without compromising defense performance. This is because the defense strength is determined by the degree of obfuscation in the mapping relationship of the soft labels. On the other hand, the mapping between original labels and soft labels is controlled by the label holder, making it challenging for attackers to access or disrupt this process. Additionally, this approach has several highlights stemming from its own design. By training with the additional features and soft labels, the approach adds a layer of protection to reduce the direct information about real labels that could be inferred from the model parameters. Finally, such a scheme is practical and straightforward to implement in real-world scenarios.
@@
  • Validate the security of existing defense works: We propose a method called the Embedding Extension Attack, which adds perturbation dimensions to the embedding to help the label inference attack[12] bypass current defense solutions.
  • Propose a robust defense mechanism against interference: We propose a method called LabObf, which maps original labels to a newly obfuscated distribution, reducing the risk of label information leakage at the cut layer.
  • Conduct detailed empirical validation: We conduct experiments on several popular datasets including Epsilon[18], Bank[19], CovType[20], and Fault Type[21]. Among these, Epsilon and Bank are binary classification datasets, while CovType and Fault Type are multi-class classification datasets. The experimental results demonstrate that our proposed approach exhibits strong defense capabilities. LabObf can significantly lower the attack success rate to below that of scenarios where the bottom model does not leak any information, while maintaining the primary task accuracy within an acceptable range.
In the following sections, we describe VFL and the existing background knowledge of attacks and defenses in Sec.2. The threat model is presented in Sec.3. We first present an attack to demonstrate the flaw in the existing defense scheme in Sec.4, and then propose a new, more robust protection scheme in Sec.5. Sec.6 and Sec.7 present experimental data and analysis, while Sec.8 analyzes related works. Finally, we conclude with a brief summary in Sec.9.
@@ 2.1 SplitNN-Based VFL
Although SplitNN can be applied to settings with multiple parties, in practice, VFL is often limited to two parties[10, 12, 14, 17, 23] due to issues of trust and efficiency. Therefore, this paper focuses exclusively on scenarios with two-party participation.
@@ 3.1 Network Architecture
Due to the widespread application of VFL tasks in practical environments involving two-party participation[10, 12, 14, 17, 23], this paper primarily focuses on the SplitNN architecture involving two participating parties, namely the host and the client, like Fig.1. In this architecture, the client, potentially acting as an attacker, holds partial dimensional features and maintains the corresponding bottom model locally. During training and prediction, it is responsible for uploading locally computed embeddings to the host. The host, acting as the defending party, holds partial dimensional features and all labels, and maintains both the corresponding bottom and top models locally. During training and prediction, it calculates the corresponding embeddings locally and combines them with the embeddings uploaded by the client, and then inputs them into the top model. The output of the top model represents the prediction result.
@@ 4 Embedding Extension Attack against Discorloss
Figure 3: The Pearson correlation between the embedding dimensions and the real labels, where $P01$ to $P04$ represent the additional perturbation dimensions.
@@ 5 Label Obfuscation Protection Scheme
The selection of soft labels significantly impacts the effectiveness of LabObf. This can be explained in three aspects: the magnitude of soft labels, the degree of cross-over between soft labels, and the number of soft labels. We analyze these three aspects and, based on experimental experience, propose the following design principles:
  1. 1.
    The magnitude of soft labels is significant. When decomposing the original label into soft labels, larger soft label values lead to larger gradients within the loss function, consequently slowing down the model convergence. Additionally, larger label values may introduce numerical instability during training. Conversely, decreasing the label values might cause the model to overlook subtle differences in the data, potentially leading to the model’s inability to accurately capture important patterns and features, thus reducing its predictive capability. Our practical experience reveals that normalizing the range of soft label values $R^{\prime}$ to match the range of original label values $R$ is a stable method that can alleviate the training difficulties introduced by label decomposition.
  2. 2.
    The degree of cross-over between soft labels is closely related to the strength of the defense. We can roughly control the defensive effect by changing the distribution of soft labels. If the soft labels of different categories do not overlap at all, for example, $0\to[0.0,0.4]$ and $1\to[0.6,1.0]$, our defense scheme would degrade to having almost no defensive effect. Conversely, if the soft labels of different categories are interleaved, the defense effect will be more ideal. Therefore, we provide an easy-to-follow design guideline for reference, i.e., for any set of $N$ soft labels randomly selected from the interval $R^{\prime}$ and ordered in ascending sequence as $\{l_{1},l_{2},\cdots,l_{N}\}$, stable and effective defense can be achieved if the following two conditions are met:
    • Intervals between soft labels should not be too small, i.e., they must satisfy: $|l_{i}-l_{j}|\geq\frac{|R^{\prime}|}{2N},\forall i,j\in\{1\ldots N\}$. This is to prevent soft labels from being too similar to each other, which would increase the difficulty of training.
    • Soft labels from the same original label should not be adjacent, i.e., they must satisfy: $Org(l_{i})\neq Org(l_{i+1}),\forall i\in\{1\ldots N-1\}$. Here, $Org(\cdot)$ denotes the function mapping a soft label to its corresponding original label. This ensures that soft labels from different original labels are appropriately interleaved.
  3. 3.
    The number of soft labels into which each original label is split also plays a crucial role in the defense’s effectiveness. Having too many soft labels may increase the learning task’s difficulty, requiring more complex models and additional iterations, while having too few soft labels may fail to achieve deep cross-interactions. Additionally, the number of soft labels affects the complexity and computational overhead of the defense scheme. Based on our practical experience, splitting each original label into $2$ soft labels is an appropriate choice. Further detailed analysis and experiments can be found in Sec.7.2.
(a) Raw model.
(b) The client has no additional attributes.
(c) The client has an additional attribute.
Figure 7: An example of how to build the mapping relationship between original classes and soft labels using the additional attributes.
@@ 6.3 Evaluation of LabObf
We validate the performance of LabObf on four datasets. In the experiments, we add one-dimensional special features, $x_{1}$ and $x_{2}$, to the attacker and the defender respectively. These features have values randomly selected from the interval $[0,200]$. Utilizing these two special features, we map each original label $l$ to a pair of soft labels $[l_{1},l_{2}]$. The specific mapping rule is as follows: when $x_{1}+x_{2}\leq 200$, $l$ is mapped to $l_{1}$; otherwise, $l$ is mapped to $l_{2}$. The selection of soft labels adheres to the design principles outlined in Sec.5. We train VFL models based on these soft labels and then applied label inference attacks to exploit predictions from the perturbed bottom model outputs.
@@ 7.1 Effect of Auxiliary Set Size on LabObf
Figure 12: Model and attack performance on binary classification tasks under different soft label numbers.
7.2 Effect of Soft Label Number on LabObf
We present a detailed analysis of how the number of soft labels influences LabObf in this section. To simplify the analysis, we assume each original label maps to an equal number of soft labels, considering three settings where each original label maps to $2$, $3$, or $4$ soft labels. Each setting is designed according to the principles for soft label magnitude and degree of cross-over in Sec.5. Given that the same settings on multiple classification datasets would generate too many soft labels, we experiment on the binary classification datasets Epsilon and Bank.
Fig.12 illustrates the model performance and attack performance under different number of soft labels. We observe that as the number of soft labels increases, the primary task accuracy decreases. This is reasonable, as an increased number of soft labels makes it harder for the model to learn and fit the soft label distribution, leading to a decrease in soft label classification accuracy, which in turn lowers the original label classification accuracy. Meanwhile, we observe that the attack success rate remains largely unchanged as the number of soft labels increases. This indicates that the model’s defense performance is more influenced by the distribution of soft label values rather than their quantity. Based on these experimental results and design principles, we find that mapping each original label to $2$ soft labels allows LabObf to achieve a satisfactory balance of primary task accuracy and label privacy protection.